Can someone help me on this problem?
I need to evaluate this expressions if a= -2, b=5, c= -4, and d=6
4b + 5d - c/a
Thanks,
Sher

Answer by awhalen(4)

4(5) + 5(6)- (-4)/(-2)=
20 + 30 - 2= 48